# Cosmological constraints from the EFT power spectrum and tree-level bispectrum of 21cm intensity maps

### v1

@@ 2.1 HI IM bias
$M_{\rm HI}(M,z)=C(z)(1-Y_{p})\frac{\Omega_{\rm b}}{\Omega_{\rm m}}\,{\rm e}^{-M_{\rm min}(z)/M}\,M^{q(z)}.$ (2.2)
@@
$\delta_{\rm HI}(\bm{x},\tau)=b_{1}(\tau)\delta_{\rm m}(\bm{x},\tau)+\varepsilon(\bm{x},\tau)+\frac{b_{2}(\tau)}{2}\delta_{\rm m}(\bm{x},\tau)^{2}+\frac{b_{s^{2}}(\tau)}{2}s(\bm{x},\tau)^{2}+\varepsilon_{\delta}(\bm{x},\tau)\,\delta_{\rm m}(\bm{x},\tau),$ (2.3)
where we dropped label ‘HI’ from $b_{i}$ for brevity. Here $\delta_{\rm m}$ is the matter overdensity and the $\varepsilon$’s are stochastic terms (see [17] for details). The first-order and quadratic biases are computed from Equation 2.1. In the redshift range $0.7\lesssim z\lesssim 2.6$, we find the fitting formulas
$b_{1}(z)$ $=1.333\leavevmode\nobreak\ \,+0.1809\;z+0.05302\;z^{2}-0.0008822\;z^{3}\;,$ (2.4)
$b_{2}(z)$ $=0.1802+0.1711\;z+0.02161\;z^{2}+0.03136\;z^{3}\;.$ (2.5)
The tidal field bias is modelled as $b_{s^{2}}=-4(b_{1}-1)/7$, following [31].
@@ 2.2 HI IM power spectrum
The HI temperature contrast is $\delta_{\rm HI}={\delta T_{\rm HI}}/{\bar{T}_{\rm HI}}$, where $T_{\rm HI}$ is the HI brightness temperature and its background value is modelled as [32]
$\bar{T}_{\rm HI}(z)=188\,h\,\Omega_{\rm HI}(z)\frac{H_{0}}{H(z)}\leavevmode\nobreak\ \mu{\rm K}\quad\mbox{with}\quad\Omega_{\rm HI}(z)=4\times 10^{-4}(1+z)^{0.6}\,.$ (2.6)
@@
$P^{11}(\bm{k},z)=Z_{1}(\bm{k},z)^{2}P_{\rm m}^{\rm L}(k,z)\quad\mbox{with}$ (2.8)
$Z_{1}(\bm{k},z)=b_{1}(z)+f(z)\mu^{2}\,.$ (2.9)
@@
$+N_{0}+N_{2}(\mu k)^{2}+N_{4}(\mu k)^{4}\,,$ (2.11)
where we substitute $P(\bm{k},z)$ with $P^{\rm EFT}(\bm{k},z)$ in Equation 2.7 and $P_{\rm SN}$ is replaced by the $N_{0}$ term. Moreover, $\alpha_{a}$ represent the counter-terms, while $N_{2}$ and $N_{4}$ encapsulate small-scale velocities. Following [34], the Zeldovich approximation of the matter power spectrum to linear order is defined as
$P^{\text{Zel}}_{\text{cb}}(\bm{k},z)=(1+f\mu^{2})^{2}P_{\rm m}^{\rm L}(k,z)\,.$ (2.12)
@@ 2.3 HI IM bispectrum
$\leavevmode\nobreak\ +2P_{\varepsilon\varepsilon_{\delta}}(z)\Big{[}Z_{1}(\bm{k}_{1},z)P_{\rm m}^{\rm L}(k_{1},z)+2\leavevmode\nobreak\ \text{perm}\Big{]}+B_{\varepsilon}(z)\bigg{\}},$ (2.13)
@@
$\leavevmode\nobreak\ +\frac{1}{2}f\mu_{ij}k_{ij}\left[\frac{\mu_{i}}{k_{i}}Z_{1}(\bm{k}_{j})+\frac{\mu_{j}}{k_{j}}Z_{1}(\bm{k}_{i})\right],$ (2.14)
@@
$D_{\text{fog}}^{B}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3},z)$ $=\exp\Big{[}-\big{(}k_{1}^{2}\mu_{1}^{2}+k_{2}^{2}\mu_{2}^{2}+k_{3}^{2}\mu_{3}^{2}\big{)}\sigma_{B}(z)^{2}\Big{]},$ (2.15)
where $\sigma_{B}$ characterizes the damping scale, with fiducial value given by the linear velocity dispersion $\sigma_{\upsilon}$. Note that we do not apply an ‘fog’ damping factor to the power spectrum since nonlinear RSD are treated by the EFT counter terms. Finally, the fiducial values of the stochastic terms in Equation 2.13 are [18]:
$P_{\varepsilon\varepsilon_{\delta}}=\frac{b_{1}}{2\overline{n}_{\rm eff}},\quad B_{\varepsilon}=\frac{1}{\overline{n}_{\rm eff}^{2}},$ (2.16)
@@ 2.4 Alcock-Paczynski effect
$P_{\rm HI}^{\rm obs}\left(k,\mu,z\right)$ $=\left(\frac{H_{\rm true}}{H_{\rm fid}}\right)\left(\frac{D_{A,\rm fid}}{D_{A,\rm true}}\right)^{2}P_{\rm HI}\left(q,\nu,z\right)\,.$ (2.17)
@@
$q(k,\mu)$ $=\beta(\mu)\,k\;,\quad\nu(k,\mu)=\frac{\mu}{\beta(\mu)}\,\frac{H_{\rm true}}{H_{\rm fid}}\quad\mbox{where}$ (2.18)
$\beta(\mu)^{2}$ $=\left(\frac{H_{\rm true}}{H_{\rm fid}}\right)^{2}\mu^{2}+\left(\frac{D_{A,\rm fid}}{D_{A,\rm true}}\right)^{2}\left(1-\mu^{2}\right)\,.$ (2.19)
Extending Equation 2.18 and Equation 2.19 to 3 dimensions, we find that [39]:
$B^{\rm obs}_{\rm HI}\left(k_{1},k_{2},k_{3},\mu_{1},\phi,z\right)$ $=\left(\frac{H_{\rm true}}{H_{\rm fid}}\right)^{2}\left(\frac{D_{A,\rm fid}}{D_{A,\rm true}}\right)^{4}B_{\rm HI}\left(q_{1},q_{2},q_{3},\nu_{1},\phi,z\right).$ (2.20)
@@ 2.5 HI IM window
$k_{\parallel}\geq A_{\rm wedge}\,k_{\perp}\quad\text{where}\quad A_{\rm wedge}=\frac{\chi(z)\,H(z)}{c(1+z)}\sin\big{[}0.61\,N_{\rm w}\,\theta_{\rm b}(z)\big{]}\,.$ (2.21)
@@
Figure 2: The smallest scale (largest $k$) assumed to be accessible as a function of redshift, using Equation 2.24 and Equation 2.25 for the power spectrum and Equation 2.22 for the bispectrum.
@@
$k_{\rm max,0}=0.25h\leavevmode\nobreak\ \mbox{(pessimistic)},\leavevmode\nobreak\ \leavevmode\nobreak\ \,0.3h\leavevmode\nobreak\ \mbox{(optimistic)}\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm Mpc}^{-1}\,.$ (2.25)
@@ 3 Fisher forecast formalism
$F_{\alpha\beta}^{P}(z_{i})=\frac{1}{2}\sum_{k}\displaystyle\int_{-1}^{1}{\rm d}\mu\,\frac{\partial P_{\rm HI}^{\rm obs}(\bm{k},_{i})}{\partial\theta_{\alpha}}\frac{\partial P_{\rm HI}^{\rm obs}(\bm{k},_{i})}{\partial\theta_{\beta}}\frac{1}{\Delta P^{2}(\bm{k},_{i})}\,$ (3.1)
@@
$F_{\alpha\beta}^{B}(z_{i})=\frac{1}{4\pi}\!\sum_{k_{1},k_{2},k_{3}}\!\int_{-1}^{1}\!\!{\rm d}\mu_{1}\!\!\int_{0}^{2\pi}\!\!{\rm d}\phi\frac{\partial B_{\rm HI}^{\rm obs}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3},z_{i})}{\partial\theta_{\alpha}}\frac{\partial B_{\rm HI}^{\rm obs}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3},z_{i})}{\partial\theta_{\beta}}\frac{1}{\Delta B^{2}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3},z_{i})},$ (3.2)
@@
$\Delta P^{2}(\bm{k},z)$ $=\frac{4\pi^{2}}{V(z)\,k^{2}\,\Delta k(z)}\,P_{\text{HI}}^{\text{obs}}(\bm{k},z)^{2}\,,$ (3.3)
$\Delta B^{2}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3},z)$ $=s_{123}\,\pi\,k_{\text{f}}(z)^{3}\,\frac{P_{\text{HI}}^{\text{obs}}(\bm{k}_{1},z)\,P_{\text{HI}}^{\text{obs}}(\bm{k}_{2},z)\,P_{\text{HI}}^{\text{obs}}(\bm{k}_{3},z)}{k_{1}\,k_{2}\,k_{3}\,[\Delta k(z)]^{3}}\,,$ (3.4)
@@
We consider a HI IM survey in interferometer mode, similar to the one proposed for HIRAX [51, 52, 53, 25] with specifications given in Table 1. The instrumental noise term that appears in the power spectrum of Equation 2.7, is assumed to be Gaussian and it is given in the case of an interferometer by [54, 55, 56]:
$P_{\rm N}(\bm{k}_{\perp},z)=T_{\rm sys}(z)^{2}\chi(z)^{2}\lambda(z)\frac{(1+z)}{H(z)}\left[\frac{\lambda(z)^{2}}{A_{e}}\right]^{2}\,\frac{S_{\rm area}}{\theta_{\rm b}(z)^{2}}\,\frac{1}{N_{\rm pol}\,n_{\rm b}(\bm{k}_{\perp},z)\,t_{\rm survey}}\,,$ (3.5)
@@
$\bm{\theta}(z_{i})=$$\Big{\{}\Omega_{\rm b},\Omega_{\rm c},h,n_{\text{s}},A_{\rm s};D_{A}(z_{i}),H(z_{i}),f(z_{i}),b_{1}(z_{i}),b_{2}(z_{i}),b_{s^{2}}(z_{i})\Big{\}}\,,$ (3.6)
@@
$F_{\alpha\beta}^{P+B}(z_{i})=F_{\alpha\beta}^{P}(z_{i})+F_{\alpha\beta}^{B}(z_{i})\,.$ (3.7)
@@
$F_{\alpha\beta}^{P+B,{\rm tot}}=\sum_{i}F_{\alpha\beta}^{P+B}(z_{i})\,,$ (3.8)
@@
$\Lambda$CDM :$:\big{\{}{\Omega_{\text{b}}},{\Omega_{\text{c}}},h,n_{\text{s}},A_{\rm s}\big{\}},$ (3.9)
$w_{0}w_{a}{\rm CDM}\,$ :$:\big{\{}{\Omega_{\text{b}}},{\Omega_{\text{c}}},h,n_{\text{s}},A_{\rm s},w_{0},w_{a}\big{\}}\quad\mbox{where}\leavevmode\nobreak\ \leavevmode\nobreak\ w_{\rm de}(z)=w_{0}+w_{a}\,\frac{z}{1+z}\,,$ (3.10)
$\gamma{\rm CDM}\,$ :$:\big{\{}{\Omega_{\text{b}}},{\Omega_{\text{c}}},h,n_{\text{s}},A_{\rm s},\gamma\big{\}}\quad\mbox{where}\leavevmode\nobreak\ \leavevmode\nobreak\ f(z)=\big{[}\Omega_{\rm m}(z)\big{]}^{\gamma}\,.$ (3.11)
In order to find the final constraints on the above parameters we apply Jacobian transformations to the Fisher matrix that corresponds to the parameter space in Equation 3.6, after marginalising over the remaining bias parameters:
$\widetilde{F}_{AB}=\big{(}J^{\rm T}\,F^{\rm tot}\,J\big{)}_{AB}=\sum_{\alpha,\beta}J_{A\alpha}\,F_{\alpha\beta}^{\rm tot}\,J_{\beta B}\qquad\mbox{where}\quad J_{A\alpha}={\partial\theta_{\alpha}\over\partial\tilde{\theta}_{A}}\,.$ (3.12)

### v2

@@ Abstract
keywords:
Cosmology , 21 cm , Fisher matrix , Effective Field Theory
journal: Physics of the Dark Universe
[…]
@@ 2.1 HI IM bias
$M_{\rm HI}(M,z)=C(z)(1-Y_{p})\frac{\Omega_{\rm b}}{\Omega_{\rm m}}\,{\rm e}^{-M_{\rm min}(z)/M}\,M^{q(z)}.$ (2)
@@
$\delta_{\rm HI}(\bm{x},\tau)=b_{1}(\tau)\delta_{\rm m}(\bm{x},\tau)+\varepsilon(\bm{x},\tau)+\frac{b_{2}(\tau)}{2}\delta_{\rm m}(\bm{x},\tau)^{2}+\frac{b_{s^{2}}(\tau)}{2}s(\bm{x},\tau)^{2}+\varepsilon_{\delta}(\bm{x},\tau)\,\delta_{\rm m}(\bm{x},\tau),$ (3)
where we dropped label ‘HI’ from $b_{i}$ for brevity. Here $\delta_{\rm m}$ is the matter overdensity and the $\varepsilon$’s are stochastic terms (see [17] for details). The first-order and quadratic biases are computed from Equation 1. In the redshift range $0.7\lesssim z\lesssim 2.6$, we find the fitting formulas
$b_{1}(z)$ $=1.333\leavevmode\nobreak\ \,+0.1809\;z+0.05302\;z^{2}-0.0008822\;z^{3}\;,$ (4)
$b_{2}(z)$ $=0.1802+0.1711\;z+0.02161\;z^{2}+0.03136\;z^{3}\;.$ (5)
The tidal field bias is modelled as $b_{s^{2}}=-4(b_{1}-1)/7$, following [31]. Note that we marginalise over the bias parameters in each redshift bin: the fitting formulas, based on a halo model framework, are used only to set the fiducial values of the bias parameters.
@@ 2.2 HI IM power spectrum
The HI temperature contrast is $\delta_{\rm HI}={\delta T_{\rm HI}}/{\bar{T}_{\rm HI}}$, where $T_{\rm HI}$ is the HI brightness temperature and its background value is related to the background HI number density $\Omega_{\rm HI}$ by [32]
$\bar{T}_{\rm HI}(z)=188\,h\,\Omega_{\rm HI}(z)\frac{H_{0}}{H(z)}\leavevmode\nobreak\ \mu{\rm K}\,.$ (6)
The background number density is currently poorly constrained and we use the model $\Omega_{\rm HI}(z)=4\times 10^{-4}(1+z)^{0.6}$, which is consistent with observations [21]. We follow the common practice of not marginalising over $\bar{T}_{\rm HI}$, assuming that future observations will determine $\Omega_{\rm HI}$. Given that $\bar{T}_{\rm HI}$ is degenerate with the bias parameters, this degrades the constraints on those parameters and on $A_{s}$ [15, 16].
@@
$P^{11}(\bm{k},z)=Z_{1}(\bm{k},z)^{2}P_{\rm m}^{\rm L}(k,z)\quad\mbox{with}$ (8)
$Z_{1}(\bm{k},z)=b_{1}(z)+f(z)\mu^{2}\,.$ (9)
@@
$+N_{0}+N_{2}(\mu k)^{2}+N_{4}(\mu k)^{4}\,,$ (11)
where we substitute $P(\bm{k},z)$ with $P^{\rm EFT}(\bm{k},z)$ in Equation 7 and $P_{\rm SN}$ is replaced by the $N_{0}$ term. Moreover, $\alpha_{a}$ represent the counter-terms, while $N_{2}$ and $N_{4}$ encapsulate small-scale velocities. Following [34], the Zeldovich approximation of the matter power spectrum to linear order is defined as
$P^{\text{Zel}}_{\text{cb}}(\bm{k},z)=(1+f\mu^{2})^{2}P_{\rm m}^{\rm L}(k,z)\,.$ (12)
@@ 2.3 HI IM bispectrum
$\leavevmode\nobreak\ +2P_{\varepsilon\varepsilon_{\delta}}(z)\Big{[}Z_{1}(\bm{k}_{1},z)P_{\rm m}^{\rm L}(k_{1},z)+2\leavevmode\nobreak\ \text{perm}\Big{]}+B_{\varepsilon}(z)\bigg{\}},$ (13)
@@
$\leavevmode\nobreak\ +\frac{1}{2}f\mu_{ij}k_{ij}\left[\frac{\mu_{i}}{k_{i}}Z_{1}(\bm{k}_{j})+\frac{\mu_{j}}{k_{j}}Z_{1}(\bm{k}_{i})\right],$ (14)
@@
$D_{\text{fog}}^{B}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3},z)$ $=\exp\Big{[}-\big{(}k_{1}^{2}\mu_{1}^{2}+k_{2}^{2}\mu_{2}^{2}+k_{3}^{2}\mu_{3}^{2}\big{)}\sigma_{B}(z)^{2}\Big{]},$ (15)
where $\sigma_{B}$ characterizes the damping scale, with fiducial value given by the linear velocity dispersion $\sigma_{\upsilon}$. Note that we do not apply an ‘fog’ damping factor to the power spectrum since nonlinear RSD are treated by the EFT counter terms. Finally, the fiducial values of the stochastic terms in Equation 13 are [18]:
$P_{\varepsilon\varepsilon_{\delta}}=\frac{b_{1}}{2\overline{n}_{\rm eff}},\quad B_{\varepsilon}=\frac{1}{\overline{n}_{\rm eff}^{2}},$ (16)
@@ 2.4 Alcock-Paczynski effect
$P_{\rm HI}^{\rm obs}\left(k,\mu,z\right)$ $=\left(\frac{H_{\rm true}}{H_{\rm fid}}\right)\left(\frac{D_{A,\rm fid}}{D_{A,\rm true}}\right)^{2}P_{\rm HI}\left(q,\nu,z\right)\,.$ (17)
@@
$q(k,\mu)$ $=\beta(\mu)\,k\;,\quad\nu(k,\mu)=\frac{\mu}{\beta(\mu)}\,\frac{H_{\rm true}}{H_{\rm fid}}\quad\mbox{where}$ (18)
$\beta(\mu)^{2}$ $=\left(\frac{H_{\rm true}}{H_{\rm fid}}\right)^{2}\mu^{2}+\left(\frac{D_{A,\rm fid}}{D_{A,\rm true}}\right)^{2}\left(1-\mu^{2}\right)\,.$ (19)
Extending Equation 18 and Equation 19 to 3 dimensions, we find that [39]:
$B^{\rm obs}_{\rm HI}\left(k_{1},k_{2},k_{3},\mu_{1},\phi,z\right)$ $=\left(\frac{H_{\rm true}}{H_{\rm fid}}\right)^{2}\left(\frac{D_{A,\rm fid}}{D_{A,\rm true}}\right)^{4}B_{\rm HI}\left(q_{1},q_{2},q_{3},\nu_{1},\phi,z\right).$ (20)
@@ 2.5 HI IM window
$k_{\parallel}\geq A_{\rm wedge}\,k_{\perp}\quad\text{where}\quad A_{\rm wedge}=\frac{\chi(z)\,H(z)}{c(1+z)}\sin\big{[}0.61\,N_{\rm w}\,\theta_{\rm b}(z)\big{]}\,.$ (21)
@@
Figure 2: The smallest scale (largest $k$) assumed to be accessible as a function of redshift, using Equation 24 and Equation 25 for the power spectrum and Equation 22 for the bispectrum.
@@
$k_{\rm max,0}=0.25h\leavevmode\nobreak\ \mbox{(pessimistic)},\leavevmode\nobreak\ \leavevmode\nobreak\ \,0.3h\leavevmode\nobreak\ \mbox{(optimistic)}\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm Mpc}^{-1}\,.$ (25)
@@ 3 Fisher forecast formalism
$F_{\alpha\beta}^{P}(z_{i})=\frac{1}{2}\sum_{k}\displaystyle\int_{-1}^{1}{\rm d}\mu\,\frac{\partial P_{\rm HI}^{\rm obs}(\bm{k},_{i})}{\partial\theta_{\alpha}}\frac{\partial P_{\rm HI}^{\rm obs}(\bm{k},_{i})}{\partial\theta_{\beta}}\frac{1}{\Delta P^{2}(\bm{k},_{i})}\,$ (26)
@@
$F_{\alpha\beta}^{B}(z_{i})=\frac{1}{4\pi}\!\sum_{k_{1},k_{2},k_{3}}\!\int_{-1}^{1}\!\!{\rm d}\mu_{1}\!\!\int_{0}^{2\pi}\!\!{\rm d}\phi\frac{\partial B_{\rm HI}^{\rm obs}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3},z_{i})}{\partial\theta_{\alpha}}\frac{\partial B_{\rm HI}^{\rm obs}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3},z_{i})}{\partial\theta_{\beta}}\frac{1}{\Delta B^{2}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3},z_{i})},$ (27)
@@
$\Delta P^{2}(\bm{k},z)$ $=\frac{4\pi^{2}}{V(z)\,k^{2}\,\Delta k(z)}\,P_{\text{HI}}^{\text{obs}}(\bm{k},z)^{2}\,,$ (28)
$\Delta B^{2}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3},z)$ $=s_{123}\,\pi\,k_{\text{f}}(z)^{3}\,\frac{P_{\text{HI}}^{\text{obs}}(\bm{k}_{1},z)\,P_{\text{HI}}^{\text{obs}}(\bm{k}_{2},z)\,P_{\text{HI}}^{\text{obs}}(\bm{k}_{3},z)}{k_{1}\,k_{2}\,k_{3}\,[\Delta k(z)]^{3}}\,,$ (29)
@@
We consider a HI IM survey in interferometer mode, similar to the one proposed for HIRAX [51, 52, 53, 25] with specifications given in Table 1. The instrumental noise term that appears in the power spectrum of Equation 7, is assumed to be Gaussian and it is given in the case of an interferometer by [54, 55, 56]:
$P_{\rm N}(\bm{k}_{\perp},z)=T_{\rm sys}(z)^{2}\chi(z)^{2}\lambda(z)\frac{(1+z)}{H(z)}\left[\frac{\lambda(z)^{2}}{A_{e}}\right]^{2}\,\frac{S_{\rm area}}{\theta_{\rm b}(z)^{2}}\,\frac{1}{N_{\rm pol}\,n_{\rm b}(\bm{k}_{\perp},z)\,t_{\rm survey}}\,,$ (30)
@@
$\bm{\theta}(z_{i})=$$\Big{\{}\Omega_{\rm b},\Omega_{\rm c},h,n_{\text{s}},A_{\rm s};D_{A}(z_{i}),H(z_{i}),f(z_{i}),b_{1}(z_{i}),b_{2}(z_{i}),b_{s^{2}}(z_{i})\Big{\}}\,,$ (31)
@@
$F_{\alpha\beta}^{P+B}(z_{i})=F_{\alpha\beta}^{P}(z_{i})+F_{\alpha\beta}^{B}(z_{i})\,.$ (32)
@@
$F_{\alpha\beta}^{P+B,{\rm tot}}=\sum_{i}F_{\alpha\beta}^{P+B}(z_{i})\,,$ (33)
@@
$\Lambda$CDM :$:\big{\{}{\Omega_{\text{b}}},{\Omega_{\text{c}}},h,n_{\text{s}},A_{\rm s}\big{\}},$ (34)
$w_{0}w_{a}{\rm CDM}\,$ :$:\big{\{}{\Omega_{\text{b}}},{\Omega_{\text{c}}},h,n_{\text{s}},A_{\rm s},w_{0},w_{a}\big{\}}\quad\mbox{where}\leavevmode\nobreak\ \leavevmode\nobreak\ w_{\rm de}(z)=w_{0}+w_{a}\,\frac{z}{1+z}\,,$ (35)
$\gamma{\rm CDM}\,$ :$:\big{\{}{\Omega_{\text{b}}},{\Omega_{\text{c}}},h,n_{\text{s}},A_{\rm s},\gamma\big{\}}\quad\mbox{where}\leavevmode\nobreak\ \leavevmode\nobreak\ f(z)=\big{[}\Omega_{\rm m}(z)\big{]}^{\gamma}\,.$ (36)
In order to find the final constraints on the above parameters we apply Jacobian transformations to the Fisher matrix that corresponds to the parameter space in Equation 31, after marginalising over the remaining bias parameters:
$\widetilde{F}_{AB}=\big{(}J^{\rm T}\,F^{\rm tot}\,J\big{)}_{AB}=\sum_{\alpha,\beta}J_{A\alpha}\,F_{\alpha\beta}^{\rm tot}\,J_{\beta B}\qquad\mbox{where}\quad J_{A\alpha}={\partial\theta_{\alpha}\over\partial\tilde{\theta}_{A}}\,.$ (37)
@@
Figure 3: Constraints on $\Lambda$CDM cosmological parameters from the combination of the EFT power spectrum and tree-level bispectrum of a HIRAX-like survey for $k_{\max,0}=0.25h/$Mpc in Equation 24. All three cases presented here consider Planck priors (see section 3) and utilise the signal from the whole redshift range.